OpenStudy (azionne):

4(7x - 1) = 3(2x + 1) - 4(x + 6) (medal)

OpenStudy (xmissalycatx):

What you're going to want to distribute the numbers in front of the parentheses, into everything inside of the parentheses. So you would multiply 4 by everything in the parentheses then you'd get 28x -4 = ________ And do that for everything on the other side. One of those has a negative, so you will multiply the values by -4! :) Hope this helped. Do you understand?

OpenStudy (smilie2015):

4(7x - 1) = 3(2x+1) - 4(x + 6) 28x - 4 = 6x + 3 - 4x - 24 28x - 6x +4x = 3 -24+ 4 26x = -17 x = -17/26